you are taking the homework. Question 12 6 pts Ima Saver has $15786 in cash. She plans to but a new car in five years. Ima is highly certain that the car will cost $28240. Ima is shopping around for a bank at which to save for the next five years. What nominal annual interest rate must the bank offer for lma to be able to afford the car, assuming that she deposits her cash today in the account? Assume annual compounding. Also assume that the bank accounts are federally insured, so that there is no risk of loss Express your answer in percent, and round your answer to three decimal places. Do not type the % symbol 4 Previous Next»
Answer #1

Future value = Present value*(1+Rate)^ Term

28240 = 15786*(1+r)^5

(1+r)^5= 28240/15786

(1+r)^5=1.788927

r= 12.336%

Answer is 12.336